OpenStudy (anonymous):

Alexei wants to hang a mirror in his boat but the mirror and frame must not have an area larger than 19.25 square feet. The mirror is 3 feet wide and 5 feet long. Which quadratic equation represents the area of the mirror and frame combined? (The border must have equal width on each side.)

OpenStudy (anonymous):

Let x = frame Area = (w+x) x (L + x) 19.25 = (3 + x)(5 + x) 19.25 = 15 + 3x + 5x +x² x² + 8x + 15 - 19.25 = 0 x² + 8x - 4.25 = 0 ◄• ANSWER!!!

OpenStudy (anonymous):

Its weird though because thats not an option. The closest answers are 4x2 + 16x − 4.25 = 0 or 2x2 + 8x − 19.25 = 0